# Observation of the quantum equivalence principle for matter-waves

### v2

@@ I Introduction
Ever since Galileo and Newton, it was clear that the phenomenon of free-fall in a gravitational field is a paradigm of physics, and must be fully understood as part of any description of nature. Einstein went a step further and used the geometry of space-time to describe gravity and free-fall. In addition, an assertion of paramount importance in physics is that to be consistent one should be able to describe nature from different frames of reference, where once more the freely-falling frame is considered fundamental. This led to the formulation of the equivalence principle, which again deals with gravity [1, 2, 3]. The rise of quantum mechanics has not changed any of these convictions, and it remains to be seen how quantum theory adheres to them. The connection between gravity (or general relativity) and quantum theory, the two pillars of modern physics, remains one of the most important open questions in physics.
@@
Here we present a direct observation of the quantum phase of a free-falling object, confirming the quantum equivalence principle for matter-waves. Measuring gravity with quantum interferometers has been previously achieved with neutrons in the celebrated Colella-Overhauser-Werner (COW) experiment [4] (as well as by Bonse and Wroblewski for a noninertial frame [5]), and with atoms [6, 7, 8, 9, 10], but the explicit phase of a free-falling wave packet has not been directly measured. This requires a unique interferometer in which one trajectory is stationary in the lab frame (relative to Earth) and acts as a reference, while the other is in free fall.
@@
We use in the following the common definition of the equivalence principle, as Einstein originally thought of it: A state of rest in a homogeneous gravitational field is physically indistinguishable from a state of uniform acceleration in a gravity-free space [11]. Let us note that numerous possible extensions of the equivalence principle to the quantum domain (hereafter the “quantum equivalence principle”) have been proposed [12, 13, 14, 15, 16, 17, 18, 19] and here we focus on the most straightforward aspect, which is the quantum phase connecting between the freely-falling frame and the lab frame [20], termed by some the Einsteinian and Newtonian frames, respectively [21]. This phase has a characteristic dependence of $t^{3}$ on the free-fall time $t$[3, 20, 21]. As this phase dependence is based on the assumption that the inertial mass is equal to the gravitational mass, as suggested by Galileo and as confirmed experimentally with early experiments by Newton, Bessel and Eötvös and more recent experiments with improved accuracy [22, 23, 24], we name our unique interferometer the Quantum Galileo Interferometer, or QGI [25].
@@
The prediction that an accelerated object, and specifically a free-falling object acquires a $t^{3}$-phase, has been made quite some time ago by Darwin, Kennard and others [26, 27, 28, 29, 30, 31, 32, 33, 34]. While a cubic phase dependence on the interferometer time may also appear in a few standard-practice interferometers [35, 36, 37, 38, 39], observing this free-fall phase in the QGI configuration sampling these two fundamental frames, constitutes a direct test and confirmation of the phase of a free-falling object, as well as a fundamental aspect of the quantum equivalence principle for matter waves [3, 17, 20].
@@
The transformation between the wavefunction $\psi_{N}$ in a Newtonian (laboratory) frame of reference, where the acceleration of an object is due to gravity, $\ddot{z}=-g$, and the wavefunction $\psi_{E}$ in an Einsteinian frame that falls in gravity together with free-falling objects, and when at $t=0$ both have the same velocity, involves a phase (e.g., Eq. 5 in Ref. [21]) which may also be referred to as a gauge phase $\phi_{\rm gauge}$[20] [see supplementary material (SM)]
@@
This fundamental phase plays two roles: first, it is the phase difference between the same wave packet as viewed from the Einsteinian and Newtonian frames in a gravitational field (e.g., Eq. 5 in Ref. [21]); second, as explained in Ref. [20], it is the phase between a wave packet in a frame accelerated with acceleration $a=-g$ in an environment devoid of gravity, and an identical wave packet in a frame that is at rest in a gravitational field, such that the two systems are indistinguishable as the phase cannot be measured by any local operation – thus heralding the quantum equivalence principle.
Our experiment intended to measure this fundamental phase — by way of interference with a reference wave packet — follows a Gedanken (thought) experiment where at time $t=t_{0}$ the wave packet is split into two paths where in one path it is at rest in the Newtonian frame, due to an applied force that is exactly opposite to gravity, while in the other path it is in free-fall, namely, at rest in the freely-falling Einsteinian frame. To achieve recombination of the two paths, the wave packet in the freely falling path is initially launched upwards by a magnetic-gradient pulse with a velocity $v_{0}$ in the direction opposite to gravity and then falls freely for a time $2T=2v_{0}/g$ (hereafter referred to as the closing condition), after which an identical magnetic-gradient pulse erases the velocity difference between the two wave packets so that the two paths overlap in position and momentum. Fig. 1(A) describes this Gedanken experiment, where the wave packet that is at rest in the lab (Newtonian) frame is termed the “reference wave packet” while the wave packet that is free-falling in the lab (Newtonian) frame is termed the “ballistic wave packet” [40].
@@
(A) Space-time diagram illustrating the classical trajectories of the wave packets in the interferometer. A beam splitter creates a coherent superposition of two wave packets with different spin states, launching one wave packet into a freely-falling ballistic trajectory for a duration of $2T$, while the second wave packet is held stationary, serving as a reference. A second beam splitter recombines the two wave packets at the end of the ballistic trajectory. State-dependent detection (not shown) measures the population in each state, revealing the phase difference between the two trajectories. (B) A schematic of our experimental realization: The stationary trajectory is held against gravity using magnetic gradients produced by currents in microfabricated wires on the atom chip. The stationary trajectory, associated with state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ of a 87Rb atom, is positioned $\approx 113\,\rm\mu m$ below the chip surface. The ballistic trajectory, associated with state $|0\rangle\equiv|F=1,m_{F}=0\rangle$, travels upwards reaching, in this example, a splitting of about $7.5\,\rm\mu m$ [about seven times the wave packet size (SM)]. The interchanging currents in the three wires (shown in the figure) create a 2D quadrupole to realize a strong magnetic gradient with a small magnetic field, which minimizes phase noise. See the Methods section for more details.
For an intuitive understanding of the QGI in the Newtonian frame, we fix the reference wave packet center at vertical position $z=0$ so that it does not accumulate any propagation phase (beyond the free evolution of the wave packet), and calculate the phase accumulated by the ballistic wave packet. The latter accumulates no propagation phase in the freely-falling (Einsteinian) frame, and we can use the gauge phase to know its phase in the Newtonian frame. We calculate the phase in the ballistic path in two parts: First, a wave packet starting at $t=0$ in the middle of the interferometer and going forward in time, and then starting from the same point and going backward in time, with these paths ending at $t=\pm T$ and $z=0$ after a free-fall duration of $T$ [Fig. 1(A)]. It follows that the gauge phase at the center of the wave packet for the two parts, and the phase difference between the two end points read
@@
In fact, the QGI also demonstrates a quantum equivalence principle between the Newtonian and Einsteinian frames, for a quantum system in a state of spatial superposition. This is due to the complete symmetry of the QGI. An observer sitting in the Einsteinian or Newtonian frames will see exactly the same phenomenon, with the roles of the reference wave packet and the ballistic wave packet inverted. This symmetry becomes strikingly clear when one computes the actions in both frames (SM). Assuming that the inertial and gravitational masses are equal and that the acceleration we induce by an external force to form the reference trajectory is exactly opposite to the gravitational acceleration, the action of the ballistic path in the Newtonian frame and that of the reference path in the Einsteinian frame are the same, equal to $-\frac{m}{3}g^{2}T^{3}$[41].
Finally, the QGI is able to measure the quantum equivalence principle for matter waves, as it utilizes no electromagnetic waves resonant with some atomic transition. In contrast, in interferometers based on light pulses, one must take into account the electromagnetic fields, namely, introduce transformations of Maxwell equations, so that a working interferometer may be described in both frames. Consequently, in such interferometers, what is in fact verified is the equivalence principle for electromagnetic waves [42, 43] (see SM). In the QGI, no such additional theoretical structure needs to be invoked.
To conclude, it may then be understood that the QGI is able to fully test the quantum equivalence principle for matter waves due to two main reasons: First, it is a configuration in which one wave packet is at rest in an inertial frame while the other is at rest in a freely-falling frame, and second, by only utilizing the theory of de-Broglie waves it achieves the complete symmetry required for testing the equivalence principle for matter waves.
@@
(A) The relative spin population in outport 1 (state $|0\rangle$) oscillates with a distinct chirp as a function of interferometer time. The blue points represent the experimental data, with error bars derived from the standard error of the mean (SEM). The black line is a fit to the data of the form $P=P_{\text{mean}}(2T)+\frac{1}{2}V(2T)\cos\left[\phi(2T)\right]$, where $V$ is the visibility of the oscillations, $\phi$ is the phase of the oscillations and $P_{\text{mean}}$ is the mean value of the population. To obtain the fit, we first identify the upper and lower envelopes and fit them to a polynomial to get $P_{\text{mean}}$ and $V(2T)$. We then fit the data to the model where the phase $\phi(2T)$ is a third-order polynomial. During the fitting we exclude the first and last oscillation as they suffer from larger uncertainty in the values of the envelope. The data shows 13 complete oscillations, with low phase noise, resulting in an average population error (SEM) of $1.8\pm 1.5$ [%] and good visibility, starting at $V=80\,\%$ and decreasing to $V=20\,\%$ at $2T=2000\,\mu\mathrm{s}$. The decrease in visibility is mainly due to a varying overlap between the wave packets due to differences in shape evolution under the influence of curved gradients as well as imperfection in the path recombination. The figure includes $633$ experimental cycles, each $30$ s, so the graph was taken over a period of $5.3$ hours. The short-term and long-term stability of the interferometer are good due to the high stability of the chip trap, and current source. (B) CCD images of the atoms in the two outports taken with on-resonance absorption imaging, showing one and a half oscillations. The top cloud is outport 1 (state $|0\rangle$), and the bottom cloud is outport 2 (state $|1\rangle$, which is magnetically sensitive), Z represents the distance from the chip, and the heatmap shows the relative optical density. The cloud in outport 2 is focused by the final splitting pulse, increasing its optical density, and making it appear slightly brighter even when the two clouds have the same population.
@@ II Experiment
Typical atom interferometers are free-space interferometers, whereby during most of their motion, the wave packets do not experience any potential applied by the experimental setup [44]. Some interferometers do have a potential applied to the wave packets during their evolution, and these may be termed, guided interferometers [45, 46]. In our case, both methods are used simultaneously, one for each of the wave packets.
In Fig. 1(B) we present the experimental setting in which the QGI, utilizing 87Rb atoms, is conducted about $113\,\mu$m below an atom chip [47], where the current-carrying wires give rise to the magnetic field gradients which are responsible for the Stern-Gerlach (SG) type forces applied to the wave packets. The QGI is a novel type of SG interferometer (SGI), leveraging a decade of SGI experiments in which the necessary techniques were developed [48], and consequently, is now capable of demonstrating enhanced stability, larger phase accumulation, and especially the flexibility required to enable one wave packet to take a ballistic trajectory while the other is stationary.
During the free-fall time, the ballistic wave packet is in state $|0\rangle\equiv|F=1,m_{F}=0\rangle$, so that to first order in the Zeeman interaction it is not affected by the magnetic gradients. The reference wave packet is in state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and is held stationary by a “holding magnetic gradient pulse” producing a magnetic force counteracting gravity. Both states belong to the atomic ground state, whereby $F$ is the total angular momentum of an atomic state and $m_{F}$ is its projection on the quantization axis.
The splitting of the wave packet is not abrupt and requires a finite kick time $T_{kick}$ in which the reference wave packet is temporarily in state $|0\rangle$ and the ballistic wave packet is temporarily in state $|1\rangle$ allowing it to be accelerated with an average acceleration $a_{kick}=v_{0}/T_{kick}$. In addition, a finite delay time $T_{d}$ is needed between the kick pulse and the holding pulse of duration $T_{h}=2T-2T_{d}$, where 2T represents the free-fall time of the ballistic wave packet. As explained in Methods, the ballistic wave packet has a free-fall time of $2T$, while the reference wave packet is in free fall during the time $T_{kick}+T_{d}$ before and after the holding time. Using $\phi_{gauge}$ these latter parts of the reference wave packet trajectory involve a free-fall phase in the Newtonian (lab) frame of
@@
Consequently, the phase of the reference wave packet is independent of the holding time.
By using arguments similar to the ones used above to derive the previous equations, we can also calculate the phase of the ballistic wave packet during the kick time and obtain the total phase difference expected in the experiment (SM)
@@
(A) The phase difference between the two interferometer arms as a function of the free-fall duration of the upper arm. The red line represents the experimental data, where the phase is extracted from Fig. 2. To estimate the uncertainty bounds, we repeat the experiment with an increase (decrease) of the magnetic-pulse current $I_{kick}$ by $0.5\,\%$ (which is our estimated experimental precision) and use the result to plot the upper (lower) uncertainty bounds. The green points are the result of a numerical simulation based on tools we developed for calculating wave packet evolution [52], which include the effects of shape and rotation of the wave packet. The blue dashed line is the result of the analytical approximation of the experiment given in Eq. 4. The analytical model assumes a homogeneous magnetic gradient, and a set of square pulses of current so that the acceleration of each arm is a piecewise constant function. The black dashed line depicts the phase of the Gedanken experiment. The inset shows the derivative of the phase with respect to $2T$. To obtain the derivative of the phase for the numeric simulation, we fit a third-order polynomial to the set of points in the main figure and take its derivative. The derivative of the analytical approximation and the Gedanken experiment are calculated by direct differentiation. The yellow dashed lines represent 5% deviations from the $T^{3}$ phase accumulation model, by modifying the analytical prediction to a power law of $T^{3\pm 0.15}$. This is meant to show that the data is not consistent with such deviations from the cubic phase. (B) The residuals between the phase of the data and the numerical simulation, where the latter is a “blind” simulation with no free parameters or fine tuning. A difference of 2 rad over 13 oscillations spanning a phase of $\approx 80$ rad, constitutes a deviation of about 2.5%. The error bars represent the statistical noise of the phase in the experiment, which shows a relative noise of 1.3% of the phase. The red band represents the systematic uncertainties [which also appear in (A)]. The blue line shows the result of the numeric simulation after fine-tuning of the currents and the magnetic field along the z-axis. The kick current is increased by 0.15% , the idle current is changed from 0.55 mA to 0.5 mA, and $B_{z}$ from 0.33 G to 0.35 G, all within the experimental uncertainties.
@@
Next, we apply a magnetic kick which launches state $|1\rangle$ upwards. Quickly thereafter, we apply a $\pi$ pulse, so that the wave packet launched upwards is in state $|0\rangle$ and is no longer affected by the magnetic gradients, thus going into a ballistic trajectory affected only by gravity (up to second order Zeeman).
Conversely, the other wave packet, which is now in state $|1\rangle$, is sensitive to the magnetic gradient which is at this time tuned so that it exactly counteracts gravity. Consequently, this wave packet remains static and does not change position. Finally, we reverse the sequence to bring the two wave packets to have the same position and momentum for a full overlap.
@@
Beyond the achievement of two very different wave packet trajectories in space, a significant experimental challenge lies in the fact that the two trajectories are very much different in momentum, and recombining the wave packets in order to observe an interference pattern requires a good overlap in both position and momentum (namely, zeroing any difference, or in other words, time-reversing the splitting process). The time-irreversibility of the splitting process, whether it be of technical or fundamental origin, has been termed by Scully, Englert and Schwinger the “Humpty-Dumpty effect” and combating it requires meticulous control over the trajectories [49].
More so, for a good overlap the size and shape of the two wave packets have to be quite similar, and as they are exposed to different magnetic gradients having different curvature, lensing effects make such an overlap hard to achieve. Finally, also rotation of the wave packets in 3D may diminish the eventual overlap, and consequently the contrast (visibility) of the interference pattern [50].
Whereas recombination in atom interferometers based on laser pulses, even when the momentum splitting is large [51], enjoys the naturally given quantum accuracy of the photon momentum, in an SGI the momentum transfer is classical in nature. This requires a high level of control over the magnetic fields as well as a detailed simulation of the wave packet evolution when exposed to such fields and gravity [52]. The QGI indeed achieves good control over position, momentum, acceleration, as well as wave packet shape and spin.

### v3

@@ I Introduction
Ever since Galileo and Newton, it was clear that the phenomenon of free-fall in a gravitational field is a paradigm of physics, and must be fully understood as part of any description of nature. Einstein went a step further and used the geometry of space-time to describe gravity and free-fall. In addition, an assertion of paramount importance in physics is that to be consistent one should be able to describe nature from different frames of reference, where once more the freely-falling frame is considered fundamental. This led to the formulation of the equivalence principle (EP), which again deals with gravity [1, 2, 3]. The rise of quantum mechanics has not changed any of these convictions, and it remains to be seen how quantum theory adheres to them. The connection between gravity (or general relativity) and quantum theory, the two pillars of modern physics, remains one of the most important open questions in physics.
@@
Here we present a direct observation of the quantum phase of a free-falling object, confirming the quantum equivalence principle (QEP) for matter-waves. Measuring gravity with quantum interferometers has been previously achieved with neutrons in the celebrated Colella-Overhauser-Werner experiment [4, 5] (as well as by Bonse and Wroblewski for a noninertial frame [6]), and with atoms [7, 8, 9, 10, 11], but the explicit phase of a free-falling wave packet (WP) has not been directly measured. This requires a unique interferometer in which one trajectory is stationary in the lab frame (relative to Earth) and acts as a reference, while the other is in free fall.
@@
The most common definition of the EP, as Einstein originally thought of it, is: A state of rest in a homogeneous gravitational field is physically indistinguishable from a state of uniform acceleration in a gravity-free space [12]. Let us note that numerous possible extensions of the equivalence principle to the quantum domain (hereafter, QEP) have been proposed [13, 14, 15, 16, 17, 18, 19, 20] and here we focus on the most straightforward aspect, which is the quantum phase connecting between the freely-falling frame and the lab frame [21], termed by some the Einsteinian and Newtonian frames, respectively [22]. This phase has a characteristic dependence of $t^{3}$ on the free-fall time $t$[3, 21, 22]. As this phase dependence is based on the assumption that the inertial mass ($m_{i}$) is equal to the gravitational mass ($m_{g}$), as suggested by Galileo and as confirmed experimentally with early experiments by Newton, Bessel and Eötvös and more recent experiments with improved accuracy [23, 24, 25], we name our unique interferometer the Quantum Galileo Interferometer, or QGI [26]. With the latter assertion, the EP is manifested in a free-falling body under gravity as viewed from accelerated and non-accelerated frames [27], or as Einstein put it: When a person is free-falling they do not feel their weight.
@@
The prediction that an accelerated object, and specifically a free-falling object acquires a $t^{3}$-phase, has been made quite some time ago by Darwin, Kennard and others [28, 29, 30, 31, 32, 33, 34, 35, 36]. While a cubic phase dependence on the interferometer time may also appear in a few standard-practice interferometers [37, 38, 39, 40, 41], observing this free-fall phase in the QGI configuration sampling these two fundamental frames, constitutes a direct test and confirmation of the phase of a free-falling object, as well as a fundamental aspect of the QEP for matter waves [3, 18, 21].
@@
The transformation between the wavefunction $\psi_{N}$ in a Newtonian (laboratory) frame of reference, where the acceleration of an object is due to gravity, $\ddot{z}=-g$, and the wavefunction $\psi_{E}$ in an Einsteinian frame that falls in gravity (with acceleration $a=-g$), and when at $t=0$ both have the same velocity, involves a phase (e.g., Eq. 5 in Ref. [22]) which may also be referred to as a gauge phase $\phi_{\rm gauge}$[21] [see supplementary material (SM)]
@@
According to the equivalence principle, $\psi_{E}$ is the wave function of a free particle that feels no forces in the falling (Einsteinian) frame (as in addition to the force $m_{g}g$ it feels a force $m_{i}a=-m_{g}g$). It is thus exactly equivalent to a particle in the lab (Newtonian) frame if gravity would be cancelled. However, Eq. (1) implies that such a particle accumulates a phase $\phi_{\rm gauge}$ in the lab frame in the presence of gravity. Measuring this fundamental phase thus confirms an outcome of the equivalence principle in the quantum domain.
Our experiment intended to measure this fundamental phase — by way of interference with a reference WP — follows a Gedanken (thought) experiment where at time $t=t_{0}$ the WP is split into two paths where in one path it is at rest in the Newtonian frame, due to an applied force that is exactly opposite to gravity, while in the other path it is in free-fall, namely, at rest in the freely-falling Einsteinian frame. To achieve recombination of the two paths, the WP in the freely falling path is initially launched upwards by a magnetic-gradient pulse with a velocity $v_{0}$ in the direction opposite to gravity and then falls freely for a time $2T=2v_{0}/g$ (hereafter referred to as the closing condition), after which an identical magnetic-gradient pulse erases the velocity difference between the two WPs so that the two paths overlap in position and momentum. Fig. 1(A) describes this Gedanken experiment, where the WP that is at rest in the lab (Newtonian) frame is termed the “reference WP” while the WP that is free-falling in the lab (Newtonian) frame is termed the “ballistic WP” [42].
@@
(A) Space-time diagram illustrating the classical trajectories of the WPs in the interferometer. A beam splitter creates a coherent superposition of two WPs with different spin states, launching one WP into a freely-falling ballistic trajectory for a duration of $2T$, while the second WP is held stationary, serving as a reference. A second beam splitter recombines the two WPs at the end of the ballistic trajectory. State-dependent detection (not shown) measures the population in each state, revealing the phase difference between the two trajectories. (B) A schematic of our experimental realization: In the longitudinal (1D) QGI interferometer, the stationary WP is held against gravity using magnetic gradients produced by currents in microfabricated wires on the atom chip. This WP, associated with state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ of a 87Rb atom, is positioned $\approx 113\,\rm\mu m$ below the chip surface. The ballistic WP, associated with state $|0\rangle\equiv|F=1,m_{F}=0\rangle$, travels upwards reaching, in this example, a splitting of about $7.5\,\rm\mu m$ [about seven times the WP size (SM)]. Interchanging currents in the three central wires (shown in the figure) create a 2D quadrupole to realize a strong magnetic gradient with a small magnetic field, which minimizes phase noise. Finally, in order to measure the population in each spin state, another (reversed) magnetic gradient separates the two states in position so that they can be imaged by a CCD. See the Methods section for more details.
For an intuitive understanding of the QGI in the Newtonian frame, we fix the reference WP center at vertical position $z=0$ so that it does not accumulate any propagation phase (beyond the free evolution of the WP), and calculate the phase accumulated by the ballistic WP. The latter accumulates no propagation phase in the freely-falling (Einsteinian) frame, and we can use the gauge phase to know its phase in the Newtonian frame. We calculate the phase in the ballistic path in two parts: First, a WP starting at $t=0$ in the middle of the interferometer and going forward in time, and then starting from the same point and going backward in time, with these paths ending at $t=\pm T$ and $z=0$ after a free-fall duration of $T$ [Fig. 1(A)]. It follows that the gauge phase at the center of the WP for the two parts, and the phase difference between the two end points read
@@
In fact, the QGI also demonstrates a QEP between the Newtonian and Einsteinian frames, for a quantum system in a state of spatial superposition. This is due to the complete symmetry of the QGI. An observer sitting in the Einsteinian or Newtonian frames will see exactly the same phenomenon, with the roles of the reference WP and the ballistic WP inverted. This symmetry becomes strikingly clear when one computes the actions in both frames (SM). Assuming that the inertial and gravitational masses are equal and that the acceleration we induce by an external force to form the reference trajectory is exactly opposite to the gravitational acceleration, the action of the ballistic path in the Newtonian frame and that of the reference path in the Einsteinian frame are the same, equal to $-\frac{m}{3}g^{2}T^{3}$[43].
Finally, the QGI is able to measure the QEP for matter waves, as it utilizes no electromagnetic waves resonant with some atomic transition. In contrast, in interferometers based on light pulses, one must take into account the electromagnetic fields, namely, introduce transformations of Maxwell equations, so that a working interferometer may be described in both frames. Consequently, in such interferometers, what is in fact verified is the EP for electromagnetic waves [44, 45] (see SM). In the QGI, no such additional theoretical structure needs to be invoked, so that only transformations of de-Broglie waves are required to understand the QGI in all frames.
To conclude, it may then be understood that the QGI is able to observe the QEP for matter waves due to two main reasons: First, it is a configuration in which one WP is at rest in an inertial frame while the other is at rest in a freely-falling frame, and second, by only utilizing the theory of de-Broglie waves it achieves the complete symmetry required for observing the QEP for matter waves.
@@
(A) The relative spin population in outport 1 (state $|0\rangle$) oscillates with a distinct chirp as a function of interferometer time. The blue points represent the experimental data, with error bars derived from the standard error of the mean (SEM). The red line is a fit to the data of the form $P=P_{\text{mean}}(2T)+\frac{1}{2}V(2T)\cos\left[\phi(2T)\right]$, where $V$ is the visibility of the oscillations, $\phi$ is the phase of the oscillations and $P_{\text{mean}}$ is the mean value of the population. To obtain the fit, we first identify the upper and lower envelopes and fit them to a polynomial to get $P_{\text{mean}}$ and $V(2T)$. We then fit the data to the model where the phase $\phi(2T)$ is a third-order polynomial. During the fitting we exclude the first and last oscillation as they suffer from larger uncertainty in the values of the envelope. The data shows 13 complete oscillations, with low phase noise, resulting in an average population error (SEM) of $1.8\pm 1.5$ [%] and good visibility, starting at $V=80\,\%$ and decreasing to $V=20\,\%$ at $2T=2000\,\mu\mathrm{s}$. The decrease in visibility is mainly due to a varying overlap between the WPs due to differences in shape evolution under the influence of curved gradients as well as imperfection in the path recombination. The figure includes $633$ experimental cycles, each $30$ s, so the graph was taken over a period of $5.3$ hours. The short-term and long-term stability of the interferometer are good due to the high stability of the chip trap, and current source. (B) Atomic distribution (heat map of optical density) adapted from CCD images (on-resonance absorption imaging), of the atoms in the two interferometer outports [see Fig. 1(B)], showing two oscillations. The top cloud is outport 1 (state $|0\rangle$), and the bottom cloud is outport 2 (state $|1\rangle$, which is magnetically sensitive). The cloud in outport 2 is focused by the final splitting pulse, increasing its optical density, and making it appear slightly denser even when the two clouds have the same population. The vertical scale of the image is 1.1 mm.
@@ II Experiment
Typical atom interferometers are free-space interferometers, whereby during most of their motion, the WPs do not experience any potential applied by the experimental setup [46]. Some interferometers do have a potential applied to the WPs during their evolution, and these may be termed, guided interferometers [47, 48]. In our case, both methods are used simultaneously, one for each of the WPs.
In Fig. 1(B) we present the experimental setting in which the QGI, utilizing 87Rb atoms, is conducted about $113\,\mu$m below an atom chip [49], where the current-carrying wires give rise to the magnetic field gradients which are responsible for the Stern-Gerlach (SG) type forces applied to the WPs. The QGI is a novel type of SG interferometer (SGI), leveraging a decade of SGI experiments in which the necessary techniques were developed [50], and consequently, is now capable of demonstrating enhanced stability, larger phase accumulation, and especially the flexibility required to enable one WP to take a ballistic trajectory while the other is stationary.
During the free-fall time, the ballistic WP is in state $|0\rangle\equiv|F=1,m_{F}=0\rangle$, so that to first order in the Zeeman interaction it is not affected by the magnetic gradients. The reference WP is in state $|1\rangle\equiv|F=2,m_{F}=1\rangle$ and is held stationary by a “holding magnetic gradient pulse” producing a magnetic force counteracting gravity. Both states belong to the atomic ground state, whereby $F$ is the total angular momentum of an atomic state and $m_{F}$ is its projection on the quantization axis.
The splitting of the WP is not abrupt and requires a finite kick time $T_{kick}$ in which the reference WP is temporarily in state $|0\rangle$ and the ballistic WP is temporarily in state $|1\rangle$ allowing it to be accelerated with an average acceleration $a_{kick}=v_{0}/T_{kick}$. In addition, a finite delay time $T_{d}$ is needed between the kick pulse and the holding pulse of duration $T_{h}=2T-2T_{d}$, where 2T represents the free-fall time of the ballistic WP. As explained in Methods, the ballistic WP has a free-fall time of $2T$, while the reference WP is in free fall during the time $T_{kick}+T_{d}$ before and after the holding time. Using $\phi_{gauge}$ these latter parts of the reference WP trajectory involve a free-fall phase in the Newtonian (lab) frame of
@@
Consequently, the phase of the reference WP is independent of the holding time.
By using arguments similar to the ones used above to derive the previous equations, we can also calculate the phase of the ballistic WP during the kick time and obtain the total phase difference expected in the experiment (SM)
@@
(A) The phase difference between the two interferometer arms as a function of the free-fall duration of the upper arm. The red line represents the experimental data, where the phase is extracted from Fig. 2. To estimate the uncertainty bounds, we repeat the experiment with an increase (decrease) of the magnetic-pulse current $I_{kick}$ by $0.5\,\%$ (which is our estimated experimental precision) and use the result to plot the upper (lower) uncertainty bounds. The green points are the result of a numerical simulation based on tools we developed for calculating WP evolution [54], which include the effects of shape and rotation of the WP. The blue dashed line is the result of the analytical approximation of the experiment given in Eq. 4. The analytical model assumes a homogeneous magnetic gradient, and a set of square pulses of current so that the acceleration of each arm is a piecewise constant function. The black dashed line depicts the phase of the Gedanken experiment. The inset shows the derivative of the phase with respect to $2T$. To obtain the derivative of the phase for the numeric simulation, we fit a third-order polynomial to the set of points in the main figure and take its derivative. The derivative of the analytical approximation and the Gedanken experiment are calculated by direct differentiation. The yellow dashed lines represent 5% deviations from the $T^{3}$ phase accumulation model, by modifying the analytical prediction to a power law of $T^{3\pm 0.15}$. This is meant to show that the data is not consistent with such deviations from the cubic phase. (B) The residuals between the phase of the data and the numerical simulation, where the latter is a “blind” simulation with no free parameters or fine tuning. A difference of 2 rad over 13 oscillations spanning a phase of $\approx 80$ rad, constitutes a deviation of about 2.5%. The error bars represent the statistical noise of the phase in the experiment, which shows a relative noise of 1.3% of the phase. The red band represents the systematic uncertainties [which also appear in (A)]. The blue line shows the result of the numeric simulation after fine-tuning of the currents and the magnetic field along the z-axis. The kick current is increased by 0.15% , the idle current is changed from 0.55 mA to 0.5 mA, and $B_{z}$ from 0.33 G to 0.35 G, all within the experimental uncertainties.
@@
Next, we apply a magnetic kick which launches state $|1\rangle$ upwards. Quickly thereafter, we apply a $\pi$ pulse, so that the WP launched upwards is in state $|0\rangle$ and is no longer affected by the magnetic gradients, thus going into a ballistic trajectory affected only by gravity (up to second order Zeeman).
Conversely, the other WP, which is now in state $|1\rangle$, is sensitive to the magnetic gradient which is at this time tuned so that it exactly counteracts gravity. Consequently, this WP remains static and does not change position. Finally, we reverse the sequence to bring the two WPs to have the same position and momentum for a full overlap.
@@
Beyond the achievement of two very different WP trajectories in space, a significant experimental challenge lies in the fact that the two trajectories are very much different in momentum, and recombining the WPs in order to observe an interference pattern requires a good overlap in both position and momentum (namely, zeroing any difference, or in other words, time-reversing the splitting process). The time-irreversibility of the splitting process, whether it be of technical or fundamental origin, has been termed by Scully, Englert and Schwinger the “Humpty-Dumpty effect” and combating it requires meticulous control over the trajectories [51].
More so, for a good overlap the size and shape of the two WPs have to be quite similar, and as they are exposed to different magnetic gradients having different curvature, lensing effects make such an overlap hard to achieve. Finally, also rotation of the WPs in 3D may diminish the eventual overlap, and consequently the contrast (visibility) of the interference pattern [52].
Whereas recombination in atom interferometers based on laser pulses, even when the momentum splitting is large [53], enjoys the naturally given quantum accuracy of the photon momentum, in an SGI the momentum transfer is classical in nature. This requires a high level of control over the magnetic fields as well as a detailed simulation of the WP evolution when exposed to such fields and gravity [54]. The QGI indeed achieves good control over position, momentum, acceleration, as well as WP shape and spin.